# Orthogonalisability of joins of graphs

### v1

@@ 4. Equivalence and non-equivalence
Theorem 3.6: (ii) does not imply (i), and Theorem 1.1: (ii) does not imply (i). To see that the converse of the first implication in Theorem 3.6 is not true, consider the pair $(\mathbf{m},\mathbf{n})=((2^{3}),(3,1^{3}))\in\mathbb{N}^{3}\times\mathbb{N}^{4}$. We have $|\mathbf{m}|=|\mathbf{n}|$, so it follows easily using Lemma 3.7(3) that (2) is not fulfilled, and hence $(\mathbf{m},\mathbf{n})$ is not strongly suitable. On the other hand, for ${\bf r}=(3,1^{3})$, the matrices
@@
satisfy $V\in\mathcal{A}({\bf r},\mathbf{m})$, $W\in\mathcal{A}({\bf r},\mathbf{n})$ and $V^{\top}W\in\mathbb{N}^{3\times 4}$, so $(\mathbf{m},\mathbf{n})\in\mathcal{D}(3,4)\subset\widetilde{\mathcal{D}}(3,4)$, hence $(\mathbf{m},\mathbf{n})$ is always join-orthogonalisable by Remark 3.2. Moreover, the same pair $(\mathbf{m},\mathbf{n})=((2^{3}),(3,1^{3}))$ does not fulfill the last inequality in Theorem 1.1(i) since in this case the left-hand side of the inequality is equal to $5$ and the right-hand side is equal to $6$, therefore Theorem 1.1(ii) does not imply (i).
@@
Theorem 3.6: (iii) does not imply (ii). Observe (either by definition or using Proposition 3.9) that the pair $(\mathbf{m},\mathbf{n})=((2^{4}),(3,1^{5}))\in\mathbb{N}^{4}\times\mathbb{N}^{6}$ is weakly suitable. However, it is not always join-orthogonalisable. To see this, suppose to the contrary; then $(\mathbf{m},\mathbf{n})\in\mathcal{D}(4,6)$. By Remark 2.3(b), there exist vectors $\boldsymbol{\delta}$ and $\boldsymbol{\varepsilon}$ with non-negative entries such that $(\mathbf{m}-\boldsymbol{\delta},\mathbf{n}-\boldsymbol{\varepsilon})\in\widetilde{\mathcal{D}}(4,6)$. Since $|\mathbf{m}|=|\mathbf{n}|$, by Remark 2.3(c) we have $|\boldsymbol{\varepsilon}|=|\boldsymbol{\delta}|$. By Lemma 2.6 we have $\mathbf{m}=(2^{4})\leq\mathbf{m}-\boldsymbol{\delta}$, so $\boldsymbol{\delta}=\boldsymbol{0}$ and hence $\boldsymbol{\varepsilon}=\boldsymbol{0}$, so $(\mathbf{m},\mathbf{n})\in\widetilde{\mathcal{D}}(4,6)$. However, $\iota_{\max}(4,6)=4<\iota(\mathbf{n})$, which contradicts Proposition 2.8.
@@
Theorem 3.6: (iv) does not imply (iii). It is easy to see that $(\mathbf{m},\mathbf{n})=((4,1),(1,1))$ is not weakly suitable, but it satisfies property (b) in Theorem 3.4, so $(\mathbf{m},\mathbf{n})$ is sometimes join-orthogonalisable.
@@
Theorem 1.1: (iv) does not imply (iii). It is straightforward to check that $(\mathbf{m},\mathbf{n}):=((2^{5}),(3,2,1^{5}))\in\mathbb{N}^{5}\times\mathbb{N}^{7}$ satisfies (iv); we claim that this pair is not always join-orthogonalisable. To see this, suppose to the contrary that $(\mathbf{m},\mathbf{n})\in\mathcal{D}(5,7)$. By a similar argument as above, Remark 2.3(b),(c) and Lemma 2.6 imply that $(\mathbf{m},\mathbf{n})\in\widetilde{\mathcal{D}}(5,7)$, so there are 00-$1$ matrices $V$ and $W$ with equal row-sums such that the column-sums of $V$ and $W$ are $\mathbf{m}=(2^{5})$ and $\mathbf{n}=(3,2,1^{5})$, respectively, and $V^{\top}W\in\mathbb{N}^{5\times 7}$. Five columns of $W$ have sum $1$; we call such columns elementary. By permuting rows of $V$ and $W$, we may assume that $\mathbf{e}_{1}^{\top}$ is a column of $W$, and since $V^{\top}W$ has no zero entry, this implies that the first row of $V$ is full of ones. Hence the first row-sum of $W$ is $5$. Since there are only two columns of $W$ with sum more than one, at least three columns of $W$ must be equal to $\mathbf{e}_{1}^{\top}$. If any of the elementary columns of $W$ were different from $\mathbf{e}_{1}^{\top}$, then by the same argument, there would be at least three such columns. Then $W$ would contain at least six elementary columns, whereas we know it has exactly five. Therefore $W$ has five columns which are all equal to $\mathbf{e}_{1}^{\top}$ and the first row of $V$ is full of ones. Let $V^{\prime}$ be the 00-$1$ matrix with column sums $(1,1,1,1,1)$ obtained by deleting the first row of $V$, and let $W^{\prime}$ be the 00-$1$ matrix with column sums $(3,2)$ obtained by deleting the first row of $W$ and the five columns equal to $\mathbf{e}_{1}^{\top}$. Since $V^{\prime}$ and $W^{\prime}$ are compatible, it follows that $((3,2),(1^{5}))\in\widetilde{\mathcal{D}}(2,5)$, which contradicts Lemma 2.6. ∎
@@
Note that Corollary 4.2 is independent of Theorem 3.6. For example, the pair $(\mathbf{m},\mathbf{n})=((2^{4}),(3,1^{5}))\in\mathbb{N}^{4}\times\mathbb{N}^{6}$ considered in the proof of Proposition 4.1 is not always join-orthogonalisable, and this follows immediately from Corollary 4.2 since $\iota_{\max}(4,6)=4<\iota(\mathbf{n})=5$. Since $(\mathbf{m},\mathbf{n})$ is weakly suitable, we cannot draw the same conclusion from Theorem 3.6. This example may be generalised, as follows.
@@ 5. Orthogonalisability of joins of graphs
In the final section of this work we return to our main motivation of determining conditions for a pair of graphs $G$ and $H$ to have $q(G\vee H)=2$. Results that follow from Section 3 are already summarised in Theorem 1.1. The next result is a direct consequence of Theorem 1.1, Proposition 3.9 and Corollary 4.4.
@@
Note that Corollary 5.1 completely characterises the always join-orthogonalisable pairs $(\mathbf{m},\mathbf{n})$, when $k\leq 2$, i.e., $\mathbf{m}\in\mathbb{N}$ or $\mathbf{m}\in\mathbb{N}^{2}$. In the next example we explain the case $k=1$.
@@
Bjorkman et al. [10, Example 4.5] proved that $q(P_{m}\vee P_{1})=\lceil\frac{m+1}{2}\rceil$ for $m\geq 2$. In particular, $q(P_{m}\vee P_{1})=2$ if and only if $1\leq m\leq 3$. If we set $k=1$ in Corollary 5.1, the latter fact generalises as follows: If $m,\ell\in\mathbb{N}$ and $\mathbf{n}\in\mathbb{N}^{\ell}$, then
@@
In particular, when $\mathbf{n}\leq(3^{\ell})$, we have $q\left(P_{m}\vee\bigcup_{i\in[\ell]}P_{n_{i}}\right)=2$ if and only if $\ell\leq m\leq|\mathbf{n}|+2$, and thus $q\left(P_{m}\vee(\ell P_{1})\right)=2$ if and only if $\ell\leq m\leq\ell+2$.
@@
In [21, Theorem 3.4] it was shown that if $G_{i}$ and $H_{i}$ are connected graphs with $\big{|}|G_{i}|-|H_{i}|\big{|}\leq 2$, $i\in[k]$, then
@@
and the other inequality holds by symmetry. So we recover [21, Theorem 3.4] as a special case.
@@ Acknowledgements
Polona Oblak and Helena Šmigoc received funding from Slovenian Research Agency, research core funding no. P1-0222.

### v2

@@ 4. Equivalence and non-equivalence
In this section, we show that while the various implications in Theorems 1.1 and 3.6 cannot generally be reversed, they may be under certain additional hypotheses.
@@
Theorem 3.6: (ii) does not imply (i), and Theorem 1.1: (ii) does not imply (i). To see that the converse of the first implication in Theorem 3.6 is not true, consider the pair $(\mathbf{m},\mathbf{n})=((2^{3}),(3,1^{3}))\in\mathbb{N}^{3}\times\mathbb{N}^{4}$. We have $|\mathbf{m}|=|\mathbf{n}|$, so it follows easily using Lemma 3.7(iii) that (2) is not fulfilled, and hence $(\mathbf{m},\mathbf{n})$ is not strongly suitable. On the other hand, for ${\bf r}=(3,1^{3})$, the matrices
@@
satisfy $V\in\mathcal{A}({\bf r},\mathbf{m})$, $W\in\mathcal{A}({\bf r},\mathbf{n})$ and $V^{\top}W\in\mathbb{N}^{3\times 4}$, so $(\mathbf{m},\mathbf{n})\in\mathcal{D}(3,4)\subset\widetilde{\mathcal{D}}(3,4)$, hence $(\mathbf{m},\mathbf{n})$ is always join-orthogonalisable by Remark 3.2. Moreover, the same pair $(\mathbf{m},\mathbf{n})=((2^{3}),(3,1^{3}))$ does not fulfil the last inequality in Theorem 1.1(i) since in this case the left-hand side of the inequality is equal to $5$ and the right-hand side is equal to $6$, therefore Theorem 1.1(ii) does not imply (i).
@@
Theorem 3.6: (iii) does not imply (ii). Observe (either by definition or using Proposition 3.9) that the pair $(\mathbf{m},\mathbf{n})=((2^{4}),(3,1^{5}))\in\mathbb{N}^{4}\times\mathbb{N}^{6}$ is weakly suitable. However, it is not always join-orthogonalisable. To see this, suppose to the contrary; then $(\mathbf{m},\mathbf{n})\in\mathcal{D}(4,6)$. By Remark 2.3(b), there exist vectors $\boldsymbol{\delta}$ and $\boldsymbol{\varepsilon}$ with non-negative entries such that $(\mathbf{m}-\boldsymbol{\delta},\mathbf{n}-\boldsymbol{\varepsilon})\in\widetilde{\mathcal{D}}(4,6)$. Since $|\mathbf{m}|=|\mathbf{n}|$, by Remark 2.3(c) we have $|\boldsymbol{\varepsilon}|=|\boldsymbol{\delta}|$. By Lemma 2.6 we have $\mathbf{m}=(2^{4})\leq\mathbf{m}-\boldsymbol{\delta}$, so $\boldsymbol{\delta}=\boldsymbol{0}$ and hence $\boldsymbol{\varepsilon}=\boldsymbol{0}$, so $(\mathbf{m},\mathbf{n})\in\widetilde{\mathcal{D}}(4,6)$. However, $\iota_{\max}(4,6)=4<\iota(\mathbf{n})$, which contradicts Proposition 2.8. For an example, see Figure 2.
a $4P_{2}\vee(P_{3}\cup 5P_{1})$
b $4P_{2}\vee(K_{3}\cup 5P_{1})$
Figure 2. The pair $(\mathbf{m},\mathbf{n})=((2^{4}),(3,1^{5}))$ is sometimes but not always join-orthogonalisable (see the proof of Proposition 4.1). This figure shows two graphs, both sized by $(\mathbf{m},\mathbf{n})$ and differing by only one edge, where the graph in (b) has $q=2$, but the graph in (a) does not. This follows immediately from Remark 3.3.
@@
Theorem 3.6: (iv) does not imply (iii). It is easy to see that $(\mathbf{m},\mathbf{n})=((4,1),(1,1))$ is not weakly suitable, but it satisfies property (ii) in Theorem 3.4, so $(\mathbf{m},\mathbf{n})$ is sometimes join-orthogonalisable.
@@
Theorem 1.1: (iv) does not imply (iii). It is straightforward to check that $(\mathbf{m},\mathbf{n}):=((2^{5}),(3,2,1^{5}))\in\mathbb{N}^{5}\times\mathbb{N}^{7}$ satisfies (iv); we claim that this pair is not always join-orthogonalisable. To see this, suppose to the contrary that $(\mathbf{m},\mathbf{n})\in\mathcal{D}(5,7)$. By a similar argument as above, Remark 2.3(b),(c) and Lemma 2.6 imply that $(\mathbf{m},\mathbf{n})\in\widetilde{\mathcal{D}}(5,7)$, so there are 00-$1$ matrices $V$ and $W$ with equal row-sums such that the column-sums of $V$ and $W$ are $\mathbf{m}=(2^{5})$ and $\mathbf{n}=(3,2,1^{5})$, respectively, and $V^{\top}W\in\mathbb{N}^{5\times 7}$. Five columns of $W$ have sum $1$; we call such columns elementary. By permuting rows of $V$ and $W$, we may assume that $\mathbf{e}_{1}^{\top}$ is a column of $W$, and since $V^{\top}W$ has no zero entry, this implies that the first row of $V$ is full of ones. Hence, the first row-sum of $W$ is $5$. Since there are only two columns of $W$ with a sum of more than one, at least three columns of $W$ must be equal to $\mathbf{e}_{1}^{\top}$. If any of the elementary columns of $W$ were different from $\mathbf{e}_{1}^{\top}$, then by the same argument, there would be at least three such columns. Then $W$ would contain at least six elementary columns, whereas we know it has exactly five. Therefore, $W$ has five columns which are all equal to $\mathbf{e}_{1}^{\top}$ and the first row of $V$ is full of ones. Let $V^{\prime}$ be the 00-$1$ matrix with column sums $(1,1,1,1,1)$ obtained by deleting the first row of $V$, and let $W^{\prime}$ be the 00-$1$ matrix with column sums $(3,2)$ obtained by deleting the first row of $W$ and the five columns equal to $\mathbf{e}_{1}^{\top}$. Since $V^{\prime}$ and $W^{\prime}$ are compatible, it follows that $((3,2),(1^{5}))\in\widetilde{\mathcal{D}}(2,5)$, which contradicts Lemma 2.6. ∎
@@
Note that Corollary 4.2 is independent of Theorem 3.6. For example, the pair $(\mathbf{m},\mathbf{n})=((2^{4}),(3,1^{5}))\in\mathbb{N}^{4}\times\mathbb{N}^{6}$ considered in the proof of Proposition 4.1 is not always join-orthogonalisable, and this follows immediately from Corollary 4.2 since $\iota_{\max}(4,6)=4<\iota(\mathbf{n})=5$. Since $(\mathbf{m},\mathbf{n})$ is weakly suitable, we cannot draw the same conclusion from Theorem 3.6. This example may be generalised as follows.
@@ 5. Orthogonalisability of joins of graphs
In the final section of this work, we return to our main motivation of determining conditions for a pair of graphs $G$ and $H$ to have $q(G\vee H)=2$. Results that follow from Section 3 are already summarised in Theorem 1.1. The next result is a direct consequence of Theorem 1.1, Proposition 3.9 and Corollary 4.4.
@@
Note that Corollary 5.1 completely characterises the always join-orthogonalisable pairs $(\mathbf{m},\mathbf{n})$, when $k\leq 2$, i.e., $\mathbf{m}\in\mathbb{N}$ or $\mathbf{m}\in\mathbb{N}^{2}$. In the next example, we explain the case $k=1$.
@@
Bjorkman et al. [11, Example 4.5] proved that $q(P_{m}\vee P_{1})=\lceil\frac{m+1}{2}\rceil$ for $m\geq 2$. In particular, $q(P_{m}\vee P_{1})=2$ if and only if $1\leq m\leq 3$. If we set $k=1$ in Corollary 5.1, the latter fact generalises as follows: If $m,\ell\in\mathbb{N}$ and $\mathbf{n}\in\mathbb{N}^{\ell}$, then
@@
In particular, when $\mathbf{n}\leq(3^{\ell})$, we have $q\left(P_{m}\vee\bigcup_{i\in[\ell]}P_{n_{i}}\right)=2$ if and only if $\ell\leq m\leq|\mathbf{n}|+2$. All such graphs with $|\mathbf{n}|\leq 3$ are illustrated in Table 1. Therefore $q\left(P_{m}\vee(\ell P_{1})\right)=2$ if and only if $\ell\leq m\leq\ell+2$.
$|\mathbf{n}|=1$ $|\mathbf{n}|=2$ $|\mathbf{n}|=3$
$m=1$
$m=2$
$m=3$
$m=4$
$m=5$
Table 1. All graphs of the form $P_{m}\vee\cup_{j=1}^{\ell}P_{n_{j}}$, having $|\mathbf{n}|\leq 3$ and $q=2$. Note that there are no such graphs for $m\geq 6$.
@@
In [24, Theorem 3.4] it was shown that if $G_{i}$ and $H_{i}$ are connected graphs with $\big{|}|G_{i}|-|H_{i}|\big{|}\leq 2$, $i\in[k]$, then
@@
and the other inequality holds by symmetry. So we recover [24, Theorem 3.4] as a special case.
@@ Acknowledgements
Polona Oblak and Helena Šmigoc received funding from the Slovenian Research Agency, research core funding no. P1-0222. The authors wish to thank the anonymous referee for their valuable comments.